# Semi-Supervised Deep Sobolev Regression: Estimation, Variable Selection and Beyond

### v1

@@ 3.1 Deep Sobolev regressor
$\widehat{f}_{{\mathcal{D}}}^{\lambda}\in\operatorname*{arg\,min}_{f\in{\mathcal{F}}}\widehat{L}_{{\mathcal{D}}}^{\lambda}(f)=\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-Y_{i})^{2}+\lambda\|\nabla f\|_{L^{2}(\nu_{X})}^{2},$ (8)
@@
The objective functional in (8), has been investigated previously within the literature of splines, according to research by Wahba (1990); Kohler and Krzyżak (2001); Kohler, Krzyżak and Schäfer (2002); Györfi et al. (2002). However, in these studies, minimization was undertaken within the Sobolev space $H^{1}(\Omega)$ or the continuous function space $C^{1}(\Omega)$ as opposed to within a class of deep neural networks.
3.2 Semi-Supervised Deep Sobolev regressor
@@
$\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}\in\operatorname*{arg\,min}_{f\in{\mathcal{F}}}\widehat{L}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}(f)=\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-Y_{i})^{2}+\frac{\lambda}{m}\sum_{i=1}^{m}\sum_{k=1}^{d}|D_{k}f(Z_{i})|^{2},$ (9)
where the deep neural network class ${\mathcal{F}}$ satisfies ${\mathcal{F}}\subseteq W^{1,\infty}(\Omega)$. A similar idea was mentioned by Belkin, Niyogi and Sindhwani (2006) in the context of manifold learning.
@@
It is worth highlighting that the measure $\nu_{X}$ is equal to $\mu_{X}$, the formulation (9) is reduced to
$\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}\in\operatorname*{arg\,min}_{f\in{\mathcal{F}}}\widehat{L}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}(f)=\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-Y_{i})^{2}+\frac{\lambda}{n+m}\sum_{i=1}^{n+m}\sum_{k=1}^{d}|D_{k}f(X_{i})|^{2},$ (10)
@@
4 Deep Sobolev Regressor with Gradient-Norm Constraint
In this section, we provide a theoretical analysis for the deep Sobolev regressor (8). The first result, given in Lemma 4.1, is an oracle-type inequality, which provides an upper-bound for the $L^{2}(\mu_{X})$-error of the deep Sobolev regressor along with an upper-bound for the $L^{2}(\nu_{X})$-norm of its gradient. Further, we show that (8) can attain the minimax optimal convergence rate, given that the regularization parameter are chosen appropriately. We also confirm that the gradient norm of the deep Sobolev regressor can be uniformly bounded by a constant.
@@ 4 Deep Sobolev Regressor with Gradient-Norm Constraint
Assumptions 2 and 3 are standard and very mild conditions in nonparametric regression, as extensively discussed in the literature (Györfi et al., 2002; Tsybakov, 2009; Nakada and Imaizumi, 2020; Schmidt-Hieber, 2020; Farrell, Liang and Misra, 2021; Jiao et al., 2023). It is worth noting that the upper-bound $B_{0}$ of hypothesis may be arbitrarily large and does not vary with the sample size $n$. In fact, this assumption can be removed through the technique of truncation, without affecting the subsequent proof, which can be found in Bauer and Kohler (2019); Kohler and Langer (2021); Kohler, Krzyżak and Langer (2022) for details.
@@
Lemma 4.1 (Oracle inequality).
Suppose Assumptions 1 to 3 hold. Let $\widehat{f}_{{\mathcal{D}}}^{\lambda}$ be the deep Sobolev regressor defined as (8) with regularization parameter $\lambda>0$. Then it follows that for each $n\geq\operatorname{VCdim}({\mathcal{F}})$,
@@
$\lesssim\inf_{f\in{\mathcal{F}}}\Big{\{}\|f-f_{0}\|_{L^{2}(\mu_{X})}^{2}+\lambda\|\nabla f\|_{L^{2}(\nu_{X})}^{2}\Big{\}}+(B_{0}^{2}+\sigma^{2})\frac{\operatorname{VCdim}({\mathcal{F}})}{n\log^{-1}(n)},$
@@
$\lesssim\inf_{f\in{\mathcal{F}}}\Big{\{}\frac{1}{\lambda}\|f-f_{0}\|_{L^{2}(\mu_{X})}^{2}+\|\nabla f\|_{L^{2}(\nu_{X})}^{2}\Big{\}}+\frac{B_{0}^{2}+\sigma^{2}}{\lambda}\frac{\operatorname{VCdim}({\mathcal{F}})}{n\log^{-1}(n)}.$
Roughly speaking, the first inequality of Lemma 4.1 decomposes the $L^{2}(\mu_{X})$-error of the deep Sobolev regressor into three terms, namely: the approximation error, the regularization term and the generalization error. Intriguingly, from the perspective of the first two terms, we need to find a deep neural network in ${\mathcal{F}}$ that not only has an sufficiently small $L^{2}(\mu_{X})$-distance from the regression function $f_{0}$, but also has an $H^{1}(\nu_{X})$-semi-norm as small as possible.
The literature on deep learning theory has extensively investigated the approximation properties of deep neural networks (Yarotsky, 2018; Yarotsky and Zhevnerchuk, 2020; Shen, Yang and Zhang, 2019; Shen, 2020; Lu et al., 2021; Petersen and Voigtlaender, 2018; Jiao et al., 2023; Li, Tang and Yu, 2019, 2020; Duan et al., 2022; Shen et al., 2023). However, there is limited research on the approximation error analysis for neural networks with gradient norm constraints (Huang et al., 2022; Jiao, Wang and Yang, 2023). The following lemma illustrates the approximation power of deep ReQU neural networks with gradient norm constraints.
Lemma 4.2 (Approximation with gradient constraints).
Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. Set the hypothesis class as a deep ReQU neural network class ${\mathcal{F}}={\mathcal{N}}(L,S)$ with $L={\mathcal{O}}(\log N)$ and $S={\mathcal{O}}(N^{d})$. Then for each $\phi\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 1}$, there exists a neural network $f\in{\mathcal{F}}$ such that
@@
This lemma provides a novel approximation error bound of deep ReQU networks with gradient norm constraint. This highlights a fundamental difference between deep ReLU and ReQU neural networks. As presented by Huang et al. (2022); Jiao, Wang and Yang (2023), the gradient norm of deep ReLU networks goes to infinity when the approximation error diminishes. In contrast, Lemma 4.2 demonstrates that deep ReQU neural networks, under a gradient norm constraint, can approximate the target function with an arbitrarily small error.
@@
Theorem 4.3 (Convergence rates).
@@
Suppose Assumptions 1 to 3 hold. Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. Assume that $f_{0}\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 1}$. Set the hypothesis class as a deep ReQU neural network class ${\mathcal{F}}={\mathcal{N}}(L,S)$ with $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d}{d+2s}})$. Let $\widehat{f}_{{\mathcal{D}}}^{\lambda}$ be the deep Sobolev regressor defined as (8) for each $\lambda>0$. Then it follows that
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}\Big{]}$ $\leq{\mathcal{O}}(\lambda)+{\mathcal{O}}\Big{(}n^{-\frac{2s}{d+2s}}\Big{)},$
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\nabla\widehat{f}_{{\mathcal{D}}}^{\lambda}\|_{L^{2}(\nu_{X})}^{2}\Big{]}$ $\leq{\mathcal{O}}(1)+{\mathcal{O}}\Big{(}\lambda^{-1}n^{-\frac{2s}{d+2s}}\Big{)}.$
Further, setting the regularization parameter $\lambda={\mathcal{O}}(n^{-\frac{2s}{d+2s}})$ implies
[…]
Here we omit the logarithmic factors and the constant behind the big ${\mathcal{O}}$ notation is independent of $n$.
@@
Theorem 4.3 quantifies how the regularization parameter $\lambda$ balances two completing goals: data fitting and the gradient norm of the estimator, and thus provides an a priori guidance for the selection of the regularization term. When one chooses $\lambda={\mathcal{O}}(n^{-\frac{2s}{d+2s}})$, the rate of the deep Sobolev regressor $\mathcal{O}(n^{-\frac{2s}{d+2s}})$ aligns with the minimax optimal rate, as established in Stone (1982); Yang and Barron (1999); Györfi et al. (2002); Tsybakov (2009). Additionally, our theoretical findings correspond to those in nonparametric regression using deep neural networks (Bauer and Kohler, 2019; Nakada and Imaizumi, 2020; Schmidt-Hieber, 2020; Kohler and Langer, 2021; Farrell, Liang and Misra, 2021; Kohler, Krzyżak and Langer, 2022; Jiao et al., 2023). In contrast to standard empirical risk minimizers, the deep Sobolev regressor imposes a constraint on the gradient norm while simultaneously ensuring the minimax optimal convergence rate. Consequently, Sobolev regularization improves the stability and enhances the generalization abilities of deep neural networks.
@@
A similar problem has been explored by researchers within the context of splines (Kohler and Krzyżak, 2001; Kohler, Krzyżak and Schäfer, 2002; Györfi et al., 2002), where the objective functional aligns with that of the deep Sobolev regressor (8). However, in these studies, minimization was token over the Sobolev space $H^{1}(\Omega)$ or the continuous function space $C^{1}(\Omega)$ instead of a deep neural network class. The consistency in this setting was studied by Kohler and Krzyżak (2001), and the convergence rate was proven to be minimax optimal by Kohler, Krzyżak and Schäfer (2002) or Györfi et al. (2002, Theorem 21.2). It is worth noting that the rate analysis in these studies relies heavily on the theoretical properties of the spline space and cannot be generalized to our setting.
5 Simultaneous Estimation of Regression Function and its Derivative
@@
$\widehat{f}_{{\mathcal{D}}}^{\lambda}\in\operatorname*{arg\,min}_{f\in\operatorname{conv}({\mathcal{F}})}\widehat{L}_{{\mathcal{D}}}^{\lambda}(f)\quad\text{and}\quad\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}\in\operatorname*{arg\,min}_{f\in\operatorname{conv}({\mathcal{F}})}\widehat{L}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}(f).$ (11)
Notice that the functions within the convex-hull $\operatorname{conv}({\mathcal{F}})$ are also deep neural networks, which can be implemented by the parallelization of neural networks (Hashem, 1997; Gühring, Kutyniok and Petersen, 2020). Therefore, in the algorithmic implementation, solving (11) will only result in mere changes compared to solving in the original problem (8) or (9).
@@
The regression function in (1) satisfies $\Delta f_{0}\in L^{2}(\nu_{X})$ and $\nabla f_{0}\cdot\bm{n}=0$ a.e. on $\partial\Omega$, where $\bm{n}$ is the unit normal to the boundary.
Since there are no measurements available on the boundary $\partial\Omega$ or out of the domain $\Omega$, it is not possible to estimate the derivatives on the boundary accurately. Hence, to simplify the problem without loss of generality, we assume that the underlying regression $f_{0}$ has zero normal derivative on the boundary, as stated in Assumption 4. This assumption corresponds to the homogeneous Neumann boundary condition in the context of partial differential equations (Evans, 2010).
@@
Lemma 5.1.
Suppose Assumptions 1, 4 and 5 hold. Let $f^{\lambda}$ be the unique minimizer of the population risk (6). Then it follows that for each $\lambda>0$
$\|f^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}$ $\lesssim\lambda^{2}\kappa\Big{\{}\|\Delta f_{0}\|_{L^{2}(\nu_{X})}^{2}+\|\nabla f_{0}\cdot\nabla(\log q)\|_{L^{2}(\nu_{X})}^{2}\Big{\}},$
$\|\nabla(f^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}$ $\lesssim\lambda\kappa\Big{\{}\|\Delta f_{0}\|_{L^{2}(\nu_{X})}^{2}+\|\nabla f_{0}\cdot\nabla(\log q)\|_{L^{2}(\nu_{X})}^{2}\Big{\}}.$
@@
5.1 Analysis for Deep Sobolev Regressor

### v2

@@ III-A Deep Sobolev regressor
$\widehat{f}_{{\mathcal{D}}}^{\lambda}\in\operatorname*{arg\,min}_{f\in{\mathcal{F}}}\widehat{L}_{{\mathcal{D}}}^{\lambda}(f)=\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-Y_{i})^{2}\\ +\lambda\|\nabla f\|_{L^{2}(\nu_{X})}^{2},$ (11)
@@
The objective functional in (11) has been investigated previously within the literature of splines, according to research by [41, 42, 43, 1]. However, in these studies, minimization was undertaken within the Sobolev space $H^{1}(\Omega)$ or the continuous function space $C^{1}(\Omega)$ as opposed to within a class of deep neural networks.
III-B Semi-Supervised Deep Sobolev regressor
@@
$\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}\in\operatorname*{arg\,min}_{f\in{\mathcal{F}}}\widehat{L}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}(f)=\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-Y_{i})^{2}\\ +\frac{\lambda}{m}\sum_{i=1}^{m}\sum_{k=1}^{d}|D_{k}f(Z_{i})|^{2},$ (12)
where the deep neural network class ${\mathcal{F}}$ satisfies ${\mathcal{F}}\subseteq W^{1,\infty}(\Omega)$. A similar idea was mentioned by [65] in the context of manifold learning.
@@
It is worth highlighting that when the measure $\nu_{X}$ is equal to $\mu_{X}$, the formulation (12) is reduced to
$\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}\in\operatorname*{arg\,min}_{f\in{\mathcal{F}}}\widehat{L}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}(f)=\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-Y_{i})^{2}\\ +\frac{\lambda}{n+m}\sum_{i=1}^{n+m}\sum_{k=1}^{d}|D_{k}f(X_{i})|^{2},$ (13)
@@
IV Deep Sobolev Regressor with Gradient-Norm Constraint
In this section, we provide a theoretical analysis for the deep Sobolev regressor (11). The first result, given in Lemma IV.1, is an oracle-type inequality, which provides an upper-bound for the $L^{2}(\mu_{X})$-error of the deep Sobolev regressor along with an upper-bound for the $L^{2}(\nu_{X})$-norm of its gradient. Further, we show that (11) attains the minimax optimal convergence rate, given that the regularization parameter are chosen appropriately. We also confirm that the gradient norm of the deep Sobolev regressor can be uniformly bounded by a constant.
@@ IV Deep Sobolev Regressor with Gradient-Norm Constraint
Assumptions 2 and 3 are standard and very mild conditions in nonparametric regression, as extensively discussed in the literature [1, 3, 35, 36, 38, 40]. It is worth noting that the upper-bound $B_{0}$ of hypothesis may be arbitrarily large and does not vary with the sample size $n$. In fact, this assumption can be removed through the technique of truncation, without affecting the subsequent proof, which can be found in [34, 37, 39] for details.
@@
Lemma IV.1 (Oracle inequality).
Suppose Assumptions 1 to 3 hold. Let $\widehat{f}_{{\mathcal{D}}}^{\lambda}$ be the deep Sobolev regressor defined as (11) with regularization parameter $\lambda>0$. Then it follows that for each $n\geq\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))$,
@@
$\lesssim\inf_{f\in{\mathcal{F}}}\Big{\{}\|f-f_{0}\|_{L^{2}(\mu_{X})}^{2}+\lambda\|\nabla f\|_{L^{2}(\nu_{X})}^{2}\Big{\}}$
$\quad+(B_{0}^{2}+\sigma^{2})\inf_{\delta>0}\Big{\{}\frac{\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))}{n}+\delta\Big{\}},$
@@
$\lesssim\inf_{f\in{\mathcal{F}}}\Big{\{}\frac{1}{\lambda}\|f-f_{0}\|_{L^{2}(\mu_{X})}^{2}+\|\nabla f\|_{L^{2}(\nu_{X})}^{2}\Big{\}}$
$\quad+\frac{B_{0}^{2}+\sigma^{2}}{\lambda}\inf_{\delta>0}\Big{\{}\frac{\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))}{n}+\delta\Big{\}}.$
Roughly speaking, the first inequality of Lemma IV.1 decomposes the $L^{2}(\mu_{X})$-error of the deep Sobolev regressor into three terms, namely: the approximation error, the regularization term, and the generalization error. Intriguingly, from the perspective of the first two terms, we need to find a deep neural network in ${\mathcal{F}}$ that not only has an sufficiently small $L^{2}(\mu_{X})$-distance from the regression function $f_{0}$, but also has an $H^{1}(\nu_{X})$-semi-norm as small as possible.
The literature on deep learning theory has extensively investigated the approximation properties of deep neural networks [73, 74, 75, 76, 77, 78, 40, 79, 80, 81, 82]. However, there is limited research on the approximation error analysis for neural networks with gradient norm constraints [83, 84]. The following lemma illustrates the approximation power of deep ReQU neural networks with gradient norm constraints.
Lemma IV.2 (Approximation with gradient constraints).
Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. Set the hypothesis class as a deep ReQU neural network class ${\mathcal{F}}={\mathcal{N}}(L,W,S)$ with $L={\mathcal{O}}(\log N)$ and $S={\mathcal{O}}(N^{d})$. Then for each $\phi\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 1}$, there exists a neural network $f\in{\mathcal{F}}$ such that
@@
This lemma provides a novel approximation error bound of deep ReQU networks with gradient norm constraint. This highlights a fundamental difference between deep ReLU and ReQU neural networks. As presented by [83, 84], the gradient norm of deep ReLU networks goes to infinity when the approximation error diminishes. In contrast, Lemma IV.2 demonstrates that deep ReQU neural networks, under a gradient norm constraint, can approximate the target function with an arbitrarily small error.
@@
Theorem IV.3 (Convergence rates).
@@
Suppose Assumptions 1 to 3 hold. Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. Assume that $f_{0}\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 1}$. Set the hypothesis class as a deep ReQU neural network class ${\mathcal{F}}={\mathcal{N}}(L,W,S)$ with $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d}{d+2s}})$. Let $\widehat{f}_{{\mathcal{D}}}^{\lambda}$ be the deep Sobolev regressor defined as (11) for each $\lambda>0$. Then it follows that
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}\Big{]}\leq{\mathcal{O}}(\lambda)+{\mathcal{O}}\Big{(}n^{-\frac{2s}{d+2s}}\log^{3}n\Big{)},$
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\nabla\widehat{f}_{{\mathcal{D}}}^{\lambda}\|_{L^{2}(\nu_{X})}^{2}\Big{]}\leq{\mathcal{O}}(1)+{\mathcal{O}}\Big{(}\lambda^{-1}n^{-\frac{2s}{d+2s}}\log^{3}n\Big{)}.$
Further, setting $\lambda={\mathcal{O}}(n^{-\frac{2s}{d+2s}}\log^{3}n)$ implies
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}\Big{]}\leq{\mathcal{O}}\Big{(}n^{-\frac{2s}{d+2s}}\log^{2}n\Big{)},$
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\nabla\widehat{f}_{{\mathcal{D}}}^{\lambda}\|_{L^{2}(\nu_{X})}^{2}\Big{]}\leq{\mathcal{O}}(1).$
Here the constant behind the big ${\mathcal{O}}$ notation is independent of $n$.
@@
Theorem IV.3 quantifies how the regularization parameter $\lambda$ balances two completing goals: data fitting and the gradient norm of the estimator, and thus provides an a priori guidance for the selection of the regularization term. When one chooses $\lambda={\mathcal{O}}(n^{-\frac{2s}{d+2s}}\log^{3}n)$, the rate of the deep Sobolev regressor $\mathcal{O}(n^{-\frac{2s}{d+2s}}\log^{3}n)$ aligns with the minimax optimal rate up to a log-factor, as established in [85, 86, 1, 3]. Additionally, our theoretical findings correspond to those in nonparametric regression using deep neural networks [34, 35, 36, 37, 38, 39, 40]. In contrast to standard empirical risk minimizers, the deep Sobolev regressor imposes a constraint on the gradient norm while simultaneously ensuring the minimax optimal convergence rate. Consequently, Sobolev regularization improves the stability and enhances the generalization abilities of deep neural networks.
@@
A similar problem has been explored by researchers within the context of splines [42, 43, 1], where the objective functional aligns with that of the deep Sobolev regressor (11). However, in these studies, minimization was token over the Sobolev space $H^{1}(\Omega)$ or the continuous function space $C^{1}(\Omega)$ instead of a deep neural network class. The consistency in this setting was studied by [42], and the convergence rate was proven to be minimax optimal by [43] or [1, Theorem 21.2]. It is worth noting that the rate analysis in these studies relies heavily on the theoretical properties of the spline space and cannot be generalized to our setting.
V Simultaneous Estimation of Regression Function and its Derivative
@@
$\widehat{f}_{{\mathcal{D}}}^{\lambda}\in\operatorname*{arg\,min}_{f\in\operatorname{conv}({\mathcal{F}})}\widehat{L}_{{\mathcal{D}}}^{\lambda}(f),\quad\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}\in\operatorname*{arg\,min}_{f\in\operatorname{conv}({\mathcal{F}})}\widehat{L}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}(f).$ (14)
Notice that the functions within the convex-hull $\operatorname{conv}({\mathcal{F}})$ are also deep neural networks, which can be implemented by the parallelization of neural networks [87, 88]. Therefore, in the algorithmic implementation, solving (14) will only result in mere changes compared to solving in the original problem (11) or (12).
@@
The regression function in (1) satisfies $\Delta f_{0}\in L^{2}(\nu_{X})$ and $\nabla f_{0}\cdot{\mathbf{n}}=0$ a.e. on $\partial\Omega$, where ${\mathbf{n}}$ is the unit normal to the boundary.
Since there are no measurements available on the boundary $\partial\Omega$ or out of the domain $\Omega$, it is not possible to estimate the derivatives on the boundary accurately. Hence, to simplify the problem without loss of generality, we assume that the underlying regression $f_{0}$ has zero normal derivative on the boundary, as stated in Assumption 4. This assumption corresponds to the homogeneous Neumann boundary condition in the context of partial differential equations [72].
@@
A sufficient condition for Assumption 5 is that $\nabla q$ is uniformly upper bounded and $q$ has a uniform positive lower bound. In fact, this stronger assumption is mild and standard for a distribution $\nu_{X}$.
@@
Lemma V.1.
Suppose Assumptions 14 and 5 hold. Let $f^{\lambda}$ be the unique minimizer of the population risk (9). Then it follows that for each $\lambda>0$
$\|f^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}$
$\lesssim\lambda^{2}\kappa\Big{\{}\|\Delta f_{0}\|_{L^{2}(\nu_{X})}^{2}+\|\nabla f_{0}\cdot\nabla(\log q)\|_{L^{2}(\nu_{X})}^{2}\Big{\}},$
$\|\nabla(f^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}$
$\lesssim\lambda\kappa\Big{\{}\|\Delta f_{0}\|_{L^{2}(\nu_{X})}^{2}+\|\nabla f_{0}\cdot\nabla(\log q)\|_{L^{2}(\nu_{X})}^{2}\Big{\}}.$
@@
V-A Analysis for Deep Sobolev Regressor